An employee of selected from a staff of 10 to supervise a certain project by selecting a tag at random from a box containing 10 tags numbered from 1to 10. Find the formula for the probability distribution of X representing the number on the tag that is drawn. What us the probability that the distribution is less than 4
1
Expert's answer
2019-07-15T09:42:32-0400

for each X the probability will be the same: P(X)=1/10=0.1

the probability that the number is less than 4: P(X<4)=P(1)+P(2)+P(3)=0.3